OpenStudy (anonymous):

Choose the correct simplification of the expression (8x − 5)(2x^2 − 5x − 6). 16x^3 − 50x^2 − 23x − 30 16x^3 − 50x^2 − 23x + 30 16x^3 − 30x^2 − 23x + 30 16x^3 + 50x^2 − 23x + 30

OpenStudy (anonymous):

16x^3 - 50x^2 − 23x + 30

OpenStudy (anonymous):

16x^2-30x^2-23x+30

OpenStudy (anonymous):

can either of u explain

OpenStudy (anonymous):

you take 8x*2x^2=16x^3 then 8x*-5x=-40x^2 then 8x*-6=-48x -5*2x^2= -10x^2 then -5*-5x=25x then -5*-6=30

OpenStudy (anonymous):

so then we have 16x^3-40x^2-48x-10x^2+25x+30 (CLT)

OpenStudy (anonymous):

thank ya !

OpenStudy (anonymous):

yuppers :)